A review of the ledger of ABC Inc. at December 31 produces the following data for the preparation of annual adjusting entries:

1. Salaries and Wages Payable, $0. There are 8 salaried employees. 6 employees receive a salary of $1,250 each per week, and 2 employees earn $680 each per week. Employees do not work weekends. All employees worked two days after the last pay period and before December 31.
2. Unearned Rent Revenue, $443,320. The company began subleasing condos in its new building on November 1. Each tenant has to make a $5,260 security deposit that is not refundable until occupancy is ended. At December 31, the company had the following rental contracts that were paid in full for the entire term of the lease:

Term ae in months) Monthly Rent of Leases Nov. 1 Dec. 1 $3,940 $10,370 3. Prepaid Advertising, $16,560. This balance consists of payments on two advertising contracts. The contracts provide for monthly advertising in two trade magazines. The terms of the contracts are as follows: Number of Contract Date Amount magazines Issues A650 May 1 $6,960 8974Oct.19,600 12 24 The first advertisement runs in the month in which the contract is signed Notes Payable, $78,480. This balance consists of a one year, 896, note that is dated June 1 4,

Prepare the adjusting entries at Dec 31

No. Date Account Titles and Explanation Debit Credit 1. Dec. 31 2. Dec. 31 3. Dec. 31 4. Dec. 31

Here are the choice list of accounts:

Accounts Payable
Accounts Receivable
Accrued Liabilities
Accumulated Amortization - Patent
Accumulated Depreciation - Buildings
Accumulated Depreciation - Equipment
Accumulated Depreciation - Furniture and Fixtures
Accumulated Depreciation - Machinery
Accumulated Depreciation - Trucks
Accumulated Impairment Losses
Accumulated OCI
Administrative Expenses
Advances to Employees
Advertising Expense
Allowance for Doubtful Accounts
Amortization Expense
Bad Debt Expense
Buildings
Cash
Common Shares
Cost of Goods Sold
Current Portion of Notes Payable
Deferred Development Costs
Delivery Expense
Depreciation Expense
Dividends
Dividends Payable
Equipment
Fair Value - NI Investments
Fair Value - OCI Investments
Freight-in
Freight-out
Furniture and Fixtures
Impairment Loss
Income Summary
Income Tax Expense
Insurance Expense
Insurance Revenue
Interest Expense
Interest Income
Interest Payable
Interest Receivable
Inventory
Investment Income or Loss
Land
Legal Expense
Loss on Inventory Due to Decline in NRV
Miscellaneous Expense
Mortgage Payable
No Entry
Notes Payable
Notes Receivable
Office Expense
Operating Expenses
Owner's Capital
Owner's Drawings
Patents
Prepaid Advertising
Prepaid Expenses
Prepaid Insurance
Prepaid Rent
Property Tax Expense
Property Tax Payable
Purchase Discounts
Purchase Returns and Allowances
Purchases
Rent Expense
Rent Payable
Rent Receivable
Rent Revenue
Repairs and Maintenance Expense
Repair Revenue
Retained Earnings
Salaries and Wages Expense
Salaries and Wages Payable
Sales Commission Expense
Sales Commission Payable
Sales Discounts
Sales Returns and Allowances
Sales Revenue
Service Revenue
Supplies
Supplies Expense
Telephone Expense
Trucks
Unearned Rent Revenue
Unearned Repair Revenue
Unearned Revenue
Unrealized Gain or Loss - OCI
Utilities Expense
Utilities Payable
Answer #1

Salaries and wages expense A/c Dr. 3544

To Salaries and wages payable A/c 3544

(On being salaries and wages for 2 days due.)

Unearned rent revenue A/c Dr. 392993.33

To rent revenue A/c    392993.33

(On being rent revenue recognised.)

Advertising expense A/c Dr. 5840

To Prepaid advertising A/c 5840

(On being advertising expense recognised.)

Interest expense A/c Dr. 3662.4

To Interest payable A/c 3662.4

( On being interest on 8% Notes payable recognised for 7 months which are due to be paid on June 1.)

Working notes:

  1. Calculation of salaries and wages payable: 2 days salaries of 6 employees at the rate $1250/week and 2 employees at the rate $680/week = [6(1250/5)*2] + [2(680/5)*2] = 3000 + 544 = $3,544
  2. Calculation of closing unearned rent revenue: It is given that rent contracts have been paid in full. So for calculating unearned rent revenue we have to calculate, (4 months lease period revenue for 6 lease contracts started on 1st nov) + (5 months lease period revenue for 4 lease contracts started on 1st dec) = [(3940/6)*4*6] + [(10370/6)*5*4] = 15760 + 34566.67 = $50,326.67. Amount of Unearned rent revenue to be debited by Rent Revenue = Unearned Rent revenue balance - closing unearned rent revenue = 443320 - 50326.67 = $392,993.33
  3. Calculation of closing prepaid advertising balance: Prepaid advertising balance is of $16,560 and also total of both the advertising contracts is $16,560 (6960 + 9600), so it means both the contracts are prepaid completely and the prepaid account contains balance of these two contracts only. Closing prepaid balance can be calculated by adding (4 months of expense of A650 contract) and (21 months of expense of B974 contract) = [(6960/12)*4] + [(9600/24)*21] = 2320 + 8400 = $10,720. Amount of Advertising expense to be debited by Prepaid advertising = Prepaid advertising balance - closing prepaid advertising balance = 16560 - 10720 = $5,840
  4. Interest payable on note: 7 months interest on notes payable from june 1 to Dec 31 for current year at the rate 8% on $78,480 = $3662.4